# Double Landmines: Invisible Textual Backdoor Attacks based on Dual-Trigger

### v2

@@ 4.1.4 Evaluation Metrics
Building on previous work Insertsent , RIPPLEWeight , Hiddenkiller , we introduce two evaluation metrics to evaluate and compare the effects of backdoor attacks. (1) Clean Accuracy (CACC): This is defined as the accuracy of the model on the original, undisturbed test data without being attacked, poisoned, or contaminated in any way. This is used to evaluate the performance of the backdoor model on clean data. (2) Attack Success Rate (ASR): This is an important metric for measuring the effectiveness of backdoor attacks. Specifically, ASR calculates the ratio of the number of times the model outputs the attacker-specified result when encountering a test sample with an embedded trigger to the total number of attacks.
@@ 4.1.5 Implementation details
The experiments in this study were mainly completed in the following hardware and software environments. In terms of GPU, most experiments used two NVIDIA RTX A6000 graphics cards (48GB of video memory each), and a small number of experiments were conducted on a dual-card NVIDIA GeForce RTX 3090 (24GB of video memory each) platform. The CPU uses AMD Ryzen Threadripper PRO 5975WX, with 32 cores and 64 threads, and a main frequency of 3.6 GHz. The memory system is configured with 4 DDR4 memory with a total of 128GB and a frequency of 3200 MHz. The operating system is Ubuntu 22.04 LTS. The model fine-tuning stage is implemented based on the LLaMAFactory framework, which supports efficient adaptation and training optimization of large language models.
According to the conclusions drawn from previous work Insertsent , the choice of target label has almost no impact on the results of backdoor attacks. Therefore, we set the target labels for the poisoned samples of these three datasets as ”Positive”, ”Offense” and ”World” respectively.
In backdoor attacks, the poisoning rate refers to the proportion of poisoned samples (samples with embedded triggers) injected into the training dataset. Different poisoning rates will significantly affect the effect of backdoor attacks. Backdoor attacks with low poisoning rates may be more covert and have less impact on model performance, but the ASR is low. Increasing the poisoning rate will correspondingly increase the ASR, but may significantly harm the CACC. Therefore, it is crucial to comprehensively compare the backdoor attack performance under different poisoning rates. After our experimental analysis, we determined that the final poisoning rates of SST-2 and OLID were 30% and 20% respectively. In contrast, the AG’s News dataset showed unique characteristics: for the Qwen2-72B-Instruct model, we determined that its final poisoning rate was 5%; while the final poisoning rates of the LLama3-8B-Instruct and LLama3.2-3B-Instruct models were both 10%.
@@
In backdoor training, we use the adamw_torch 5-1 optimizer and set the initial learning rate to 5.0e-4 (we use the most advanced fine-tuning architecture to fine-tune the victim model, and in the process decided not to use the learning rate setting in the baseline method. In previous studies, for example, Hiddenkiller used 2e-5 as the initial learning rate; however, according to our experimental reproduction results, the use of such a learning rate has led to performance regression. Therefore, we choose to explore more efficient learning strategies to achieve the best model performance). The ratio of the learning rate warmup phase is 0.1, and the type of the learning rate scheduler is ”cosine”. The learning rate of each batch is decayed by cosine annealing, The process is formulated by (6) 5-2 . In addition, we train the three models Qwen2-72B-Instruct, LLama3-8B-Instruct, and LLama3.2-3B-Instruct for different rounds. Specifically, they are trained for 5, 4, and 3 rounds respectively. The other hyperparameters and training settings of the baseline are the same as its original implementation.
@@ 4.3 Results of the attack with only the subjunctive activated
4.4 Impact of poisoning rate
In this section, we study the impact of different poisoning rates on the performance of the dual-trigger backdoor attack. As shown in Fig. 4, as the poisoning rate increases, the ASR initially shows an upward trend; however, when the poisoning rate is very high, the ASR begins to fluctuate and even decreases. On the other hand, as the poisoning rate increases, the CACC generally shows a downward trend. These results show the trade-off between ASR and CACC in backdoor attacks.
In fact, we cannot determine the final poisoning rate of the model based on the above results alone. For example, when evaluating the Qwen2-72B-Instruct model, we set the final poisoning rates of 30%, 20%, and 5% for the SST-2, OLID, and AG’s News datasets, respectively. In addition, our research also shows that the impact of defense measures on attack performance should also be considered when selecting the poisoning rate. Therefore, determining the final poisoning rate is a process that requires comprehensive consideration of multiple factors.
In order to evaluate the impact of defense measures on the poisoning rate, this paper conducted a special study based on the AG’s News dataset. As shown in Fig. 5(a), for the Qwen2-72B-Instruct model, when the poisoning rate reached 10%, its ASR and CACC performed well; however, after adopting the Onion defense strategy, CACC dropped significantly from 89.91% at point A to 34.93% at point B. In contrast, when the poisoning rate was only 5%, CACC was relatively less affected by Onion defense, so we determined 5% as its final poisoning rate. According to Fig. 5(b) and 5(c), for the other two models, we determined 10% as their final poisoning rate.
Finally, we came to the following conclusions: For the AG’s News dataset, the final poisoning rates of the three models were set to 5%, 10%, and 10%, respectively. For the SST-2 and OLID datasets, the final poisoning rates of the three models were unified to 30% and 20%, respectively.
(a) Case 1: Qwen2-72B-Instruct.
(b) Case 2: LLama3-8B-Instruct.
(c) Case 3: LLama3.2-3B-Instruct.
Figure 5: The impact of ONION on the poisoning rate selection.
@@ 5 Attack Invisibility and Robustness against defense strategies
In this section, we will explore in depth the invisibility of different backdoor attack methods and their robustness against defense strategies. Among them, the invisibility of the trigger is a key factor in backdoor attacks, refers to the fact that the poisoned samples are indistinguishable from normal samples, making it difficult to distinguish them 6-1 . Highly invisible design can effectively bypass manual or automated data review processes, greatly reducing the risk of being detected and removed. Given that many defense strategies rely on meticulous data review, there is a close connection between the invisibility of backdoor attacks and their robustness against defense strategies. In order to evaluate the invisibility of dual-trigger, we used several automated data quality evaluation methods to evaluate the differences between the poisoned samples it generated and normal samples.
@@ 5.1 Automated Data Quality Evaluation
These three automated data quality evaluation strategies specifically include: semantic similarity analysis (SSA) based on the ”all-MiniLM-L6-v2”444”all-MiniLM-L6-v2” is a pre-trained language model provided by Microsoft, belonging to the MiniLM series. It is a small but efficient transformer model, especially suitable for tasks such as semantic text similarity, text classification, and information retrieval. language model, perplexity (PPL) evaluated using the GPT-2 model 6-2 , and comprehensive text quality checks using the LanguageTool [^5] https://www.languagetool.org. The latter is further divided into four dimensions: overall text score (TS), number of spelling errors (SEN), number of grammatical errors (GEN), and number of style issues (SIN). When using LanguageTool to evaluate text quality, for each backdoor attack method, the number of GEN is calculated based on 600 samples randomly selected from the corresponding poisoned test dataset. The other three indicators are average values calculated based on the overall performance of the entire poisoned test dataset. This comprehensive consideration method aims to provide a more comprehensive and accurate text quality evaluation system.
@@ 5.2 Robustness against defense strategies
Nowadays, people have fully realized the threat posed by backdoor attacks, and have proposed a series of defense strategies to cope with this challenge. At present, a variety of textual backdoor defense strategies have been widely used: (1) ONION ONION : It is based on test sample inspection and can be applied to any victim model. (2) Back-translation Paraphrasing: The core idea of this method is to first translate the data sample into Chinese, and then translate it back to English, hoping that the trigger factors embedded in the test sample can be eliminated through paraphrasing. (3) Syntactic Structure Alteration: This scheme relies on a specific algorithm to process each sample and reconstruct it into a version with an extremely common syntactic structure ”S(NP)(VP)(.)”. The purpose of this is to reduce or even eliminate those special syntactic features that may be used to activate hidden backdoor attacks. In this section, we will analyze and compare the resistance of different backdoor attack methods to the above defense strategies one by one.
@@ 7 CONCLUSION AND FUTURE WORK
This paper also optimizes the final poisoning rate selection strategy of the baseline method. In addition to comparing key indicators such as ASR and CACC at different poisoning rates, we also comprehensively consider multiple factors such as the interference of Onion defense strategy on attack performance to determine the most appropriate final poisoning rate. In this process, we concluded that the interference of defense measures on attack performance cannot be ignored when determining the final poisoning rate of backdoor attacks. This is a process that takes multiple factors into consideration.

### v3

@@ 4.1.4 Evaluation Metrics
Building on previous work [13], [14], [15], we introduce two evaluation metrics to evaluate and compare the effects of backdoor attacks. (1) Clean Accuracy (CACC): This is defined as the accuracy of the model on the original, undisturbed test data without being attacked, poisoned, or contaminated in any way. This is used to evaluate the performance of the backdoor model on clean data. (2) Attack Success Rate (ASR): This is an important metric for measuring the effectiveness of backdoor attacks. Specifically, ASR calculates the ratio of the number of times the model outputs the attacker-specified result when encountering a test sample with an embedded trigger to the total number of attacks.
@@ 4.1.5 Implementation details
The experiments in this study were mainly completed in the following hardware and software environment.
Computing equipment:
GPU: Most experiments use two NVIDIA RTX A6000 graphics cards (48GB video memory each), and a small number of experiments are performed on two NVIDIA GeForce RTX 3090 graphics cards (24GB video memory each).
CPU: AMD Ryzen Threadripper PRO 5975WX (32 cores 64 threads, main frequency 3.6 GHz). Memory: 128GB DDR4 memory (4 × 32GB, 3200 MHz).
Software environment:
Operating system: Ubuntu 22.04 LTS.
Fine-tuning framework: Model fine-tuning is implemented based on the LLaMAFactory framework, which supports efficient adaptation and training optimization of large language models.
According to the conclusions drawn from previous work [13], the choice of target label has almost no impact on the results of backdoor attacks. Therefore, we set the target labels for the poisoned samples of these three datasets as ”Positive”, ”Offense” and ”World” respectively.
@@
In backdoor training, we use the adamw_torch [36] optimizer and set the initial learning rate to 5.0e-4 (we use the most advanced fine-tuning architecture to fine-tune the victim model, and in the process decided not to use the learning rate setting in the baseline method. In previous studies, for example, [15] used 2e-5 as the initial learning rate; however, according to our experimental reproduction results, the use of such a learning rate has led to performance regression. Therefore, We explore using a more effective learning rate schedule). The ratio of the learning rate warmup phase is 0.1, and the type of the learning rate scheduler is ”cosine”. The learning rate of each batch is decayed by cosine annealing, The process is formulated by (6) [37]. In addition, we train the three models Qwen2-72B-Instruct, LLama3-8B-Instruct, and LLama3.2-3B-Instruct for different rounds. Specifically, they are trained for 5, 4, and 3 rounds respectively. The other hyperparameters and training settings of the baseline are the same as its original implementation.
@@ 5 Attack Invisibility and Robustness against defense strategies
In this section, we will explore in depth the invisibility of different backdoor attack methods and their robustness against defense strategies. Among them, the invisibility of the trigger is a key factor in backdoor attacks, refers to the fact that the poisoned samples are indistinguishable from normal samples, making it difficult to distinguish them [38]. Highly invisible design can effectively bypass manual or automated data review processes, greatly reducing the risk of being detected and removed. Given that many defense strategies rely on meticulous data review, there is a close connection between the invisibility of backdoor attacks and their robustness against defense strategies. In order to evaluate the invisibility of dual-trigger, we used several automated data quality evaluation methods to evaluate the differences between the poisoned samples it generated and normal samples.
@@ 5.1 Automated Data Quality Evaluation
These three automated data quality evaluation strategies specifically include: semantic similarity analysis (SSA) based on the ”all-MiniLM-L6-v2”444”all-MiniLM-L6-v2” is a pre-trained language model provided by Microsoft, belonging to the MiniLM series. It is a small but efficient transformer model, especially suitable for tasks such as semantic text similarity, text classification, and information retrieval. language model, perplexity (PPL) evaluated using the GPT-2 model [39], and comprehensive text quality checks using the LanguageTool [^5] https://www.languagetool.org. The latter is further divided into four dimensions: overall text score (TS), number of spelling errors (SEN), number of grammatical errors (GEN), and number of style issues (SIN). When using LanguageTool to evaluate text quality, for each backdoor attack method, the number of GEN is calculated based on 600 samples randomly selected from the corresponding poisoned test dataset. The other three indicators are average values calculated based on the overall performance of the entire poisoned test dataset. This comprehensive consideration method aims to provide a more comprehensive and accurate text quality evaluation system.
@@ 5.2 Robustness against defense strategies
Nowadays, people have fully realized the threat posed by backdoor attacks, and have proposed a series of defense strategies to cope with this challenge. At present, a variety of textual backdoor defense strategies have been widely used: (1) ONION [26]: It is based on test sample inspection and can be applied to any victim model. (2) Back-translation Paraphrasing: The core idea of this method is to first translate the data sample into Chinese, and then translate it back to English, hoping that the trigger factors embedded in the test sample can be eliminated through paraphrasing. (3) Syntactic Structure Alteration: This scheme relies on a specific algorithm to process each sample and reconstruct it into a version with an extremely common syntactic structure ”S(NP)(VP)(.)”. The purpose of this is to reduce or even eliminate those special syntactic features that may be used to activate hidden backdoor attacks. In this section, we will analyze and compare the resistance of different backdoor attack methods to the above defense strategies one by one.
@@ 7 CONCLUSION AND FUTURE WORK
Acknowledgements
We would like to extend our heartfelt gratitude to the editor and the review-ers ofthis paper. Their professional advice and valuable feedback significantlycontributed to the quality of this research.
Author contributions
All authors read and approved the manuscript.
Funding
This work was supported by the Joint Funds of National Natural Science Foundation of China (Grant No. U23A20304).
Availablity of data and materials
Applicable.
Declarations
Competing interests
The authors declare that they have no Conflict ofinterest.